OpenStudy (anonymous):

i need help -28-(-40)=

OpenStudy (anonymous):

remove the bracket!!!

OpenStudy (anonymous):

-28-40=

Parth (parthkohli):

-x-(-y) = -x + y

OpenStudy (turingtest):

not quite reminder all: do not just give answer

OpenStudy (anonymous):

- * - = + so it would look like this -28 + 40 = so it equals 12

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

ohhh duhh

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok tell me one thing \[-1\times-1=??\]

OpenStudy (turingtest):

no need for apologies :)

OpenStudy (anonymous):

to sheg, it would result in 1

OpenStudy (anonymous):

lol alright i hate math tbh

OpenStudy (anonymous):

yes so -28 -(-40) can be written as

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

like what @myrah

OpenStudy (anonymous):

-28-(-40)